OpenStudy (anonymous):

Find the inverse of the function. f(x) = x^3 + 8

OpenStudy (anonymous):

x^3 = f(x) - 8 raise both sides to the power of (1/3) and get: x = (f(x) - 8)^(1/3) ill denote f-1(x) as the inverse of f(x) so : f-1(x) = (x - 8)^(1/3)